OpenStudy (anonymous):

there are 4 bells ringing after every 6,7,8 and 10 seconds, If they have just rung together, how many more times they ring together in next one hour ?

OpenStudy (anonymous):

Find LCM of 6,7,8 and 10=? =x s (say) no. of times it rings in one hour \[=\frac{ 3600 }{ x }=?\]

OpenStudy (anonymous):

but 3600 comes from what ? ?

OpenStudy (anonymous):

1 hour=1*60 minutes =1*60*60=3600 seconds.

OpenStudy (anonymous):

But How & from where you got the hint to assume 2 hour cause in question only next 1 hour is specified . .

OpenStudy (anonymous):

i did not say after 2 hours ,it is one hour

OpenStudy (anonymous):

then why you did =1*60*60=3600 seconds. Can You Please Elaborate . .

OpenStudy (whpalmer4):

1 hour * 60 minutes/1 hour * 60 seconds / 1 minute = 3600 seconds

OpenStudy (anonymous):

Ok Sorry sorry . . :p Got it . .

OpenStudy (anonymous):

 it is a natural number ,leave the fraction,take only natural number.

OpenStudy (anonymous):

whpalmer4 Thank You . . & Surjithayer Thnak You So Much for Brief Mention . . . :))